# On the Verification of Control Flow Attestation Evidence

### v2

@@ 4. SABRE
$loop_{count}$ a special entry in $\mathcal{C}F_{Log}$ to optimize static loops (i.e., loops with no internal branches) (Nunes et al., 2021b; Caulfield et al., 2023); denotes total times the previous $\mathcal{C}F_{Log}$ entry repeated
@@
4.1. Static Analysis and Path Verifier
Figure 4. CFG and $\mathcal{C}F_{Log}$
Figure 5. Backward Traversal for a ret
Figure 6. Backward Traversal for an call
@@ 4.2. Exploit Locator
Phase 2. Symbolic DF Analysis. Within $\mathcal{C}F_{Slice}$ determined by Phase 1, Exploit Locator must locate the instruction that corrupts $addr_{target}$, typically requiring emulation of $\mathcal{C}F_{Slice}$ execution. Without access to run-time data prior to $\mathcal{C}F_{Slice}$ execution, Exploit Locator emulates this execution using symbolic analysis. Unlike traditional symbolic execution (King, 1976), we do not use symbolic analysis to explore different paths. Instead, we use it to reconstruct all possible data flows during $\mathcal{C}F_{Slice}$ execution (Yagemann et al., 2021b, a) and monitor for the change in data flows that result in corruption of $addr_{target}$.
@@ 4.3. Patch Generator
The goal of this module is to implement a binary patch that fixes the detected vulnerability. The exact steps depend on the detected root cause (either use-after-free or buffer overflow).
Patching Use-After-Free.
@@
Patching Buffer Overflows.
@@
The example of Fig. 8 depicts the two steps required to obtain the buffer’s bounds. The previous phase has detected that the instruction mov r14, 0(r15) triggers a write to base, setting $addr_{Acc}$ to 0xe290. From this instruction, this phase knows that its destination register, r15, stores the write address pointing into the overflown buffer, and thus labels it as $reg_{Acc}$. To determine the start address of the overflown buffer, SABRE performs a backward traversal from 0xe290 until it finds the instruction that initializes the buffer’s lower bound.
@@
Although this strategy could overestimate (i.e., the estimated bounds may include other data objects and the buffer of interest), narrowing the bounds any further cannot be done without source-code level semantics. For example, the instruction at 0xe0ea modifies the stack pointer by ten. However, it cannot be determined through the binary alone whether this is allocating space for a buffer of size ten or for all stack variables totaling ten memory addresses. Although this differentiation cannot be made through binary alone, SABRE’s approach guarantees overflows into $addr_{target}$ are blocked by restricting overflows within the buffer’s own stack frame.
@@
The call instruction at 0xe106 previously called to the “unsafe” function at 0xe2fc. In the patched version, this call instruction now calls 0xe61a, the location of the patched version of the function. As depicted, this new function now contains the bounds-checking logic from T3.
@@ 4.4. Patch Validator
Figure 10. Run-time of SABRE submodules (ms) to analyze evidence of buffer overflow (OVF) and use-after-free (UAF) vulnerabilities on two MCUs and CFA architectures: hardware-based CFA atop MSP430 (Caulfield et al., 2023) and TEE-based CFA atop ARM Cortex-M33 (Caulfield et al., 2024a).
@@ 5. Implementation & Evaluation
SABRE was implemented as a combination of Bash and Python and evaluated on an Ubuntu 18.04 machine with an Intel Core i7-4790 CPU at 3.60GHz and 32GB RAM. BEEBs benchmark applications for embedded platforms (Pallister et al., 2013) were used for the evaluation. They were compiled for MSP430 and ARM Cortex-M33 and executed atop unmodified CFA architectures, ACFA (Caulfield et al., 2023) and TRACES (Caulfield et al., 2024a), to generate a dataset of executables and corresponding $\mathcal{C}F_{Log}$-s. The MSP430 prototype was configured with 8 KB of application program memory, and the ARM Cortex-M33 prototype was configured with 256 KB of application program memory. We believe that SABRE concepts generalize to several CFA architectures and use ACFA and TRACES as case studies due to their open-source availability.
To obtain CFGs from executables, SABRE first runs objdump over the executables to decode the assembly instructions. Then, it builds the CFG from this disassembled file. SABRE uses SymPy (Team, 2024) for evaluating symbolic expressions and Keystone assembler (Keystone, 2024) for generating ARM Cortex-M instructions. Since no MSP430 assembler exists in Python packages, we build upon and modify MSProbe (Swiftloke, 2024) for MSP430 instructions. Finally, SABRE uses the Python library ELFtools (Foundation, 2024c) to generate the final patched executable. SABRE proof of concept prototype is publicly available at (Caulfield et al., 2024b).
@@ 5.1. Run-times of SABRE Sub-Modules
We evaluate the run-time of each SABRE submodule after a malicious path was detected. Given the lack of reproducible exploits and incomplete CVEs for MCUs software (e.g., missing source codes, broken links, etc.) (Tan et al., 2024), we craft control flow attacks that exploit buffer overflow and use-after-free vulnerabilities, following a similar methodology to recent related work (Chilese et al., 2024). We then insert them into the example benchmark applications (Pallister et al., 2013) to closely mimic existing CVEs related to control flow hijacking and ROP in MCUs (National Vulnerability Database, 2021a, 2017, 2020a, 2020b, 2020c, b, 2022) to the best of our ability. For consistency across applications, we insert the vulnerability into the last function called by the main. The vulnerability is used to overwrite $addr_{target}$ to return/jump to the main function, causing an infinite loop. We execute the vulnerable programs on two unmodified CFA architectures – a TEE-based approach built atop ARM Cortex-M33 (TRACES (Caulfield et al., 2024a)) and a custom hardware-based approach built atop TI MSP430 (ACFA (Caulfield et al., 2023)) – to collect $\mathcal{C}F_{Log}$-s of the attack, which are used by SABRE with the modified application binary (App.elf) for further analysis. Fig. 10 shows the run-time of each SABRE submodule during the analysis.
@@
Path Verifier. Fig. 10(b) presents the time to verify $\mathcal{C}F_{Log}$. The verification time depends on the number of branch instructions executed by the program before the corrupted $addr_{target}$ is used. In most cases, the crafted OVF generates a larger $\mathcal{C}F_{Log}$ since it introduces more control flow transfers before corrupting $addr_{target}$ compared to the crafted UAF examples. Additionally, $\mathcal{C}F_{Log}$-s of ARM executables (generated by TRACES (Caulfield et al., 2024a)) require longer run-time to verify than $\mathcal{C}F_{Log}$-s of MSP430 executables (generated by ACFA (Caulfield et al., 2023)). This is because $\mathcal{C}F_{Log}$-s from TRACES do not contain any static branch instruction destinations, whereas $\mathcal{C}F_{Log}$-s from ACFA contain all control flow transfers. Consequently, TRACES $\mathcal{C}F_{Log}$-s require more time to verify since more processing is required by $\sf{\mathcal{V}rf}$ to account for the missing entries, whereas ACFA $\mathcal{C}F_{Log}$-s can be naively followed by $\sf{\mathcal{V}rf}$ without additional effort. The detection time through $\mathcal{C}F_{Log}$ verification for the evaluated examples ranged from $\approx$1.14 to 34.74ms.
@@
Similar to the Exploit Locator, the run-time of the Patch Validator is dominated by the Symbolic DF Analysis (from 0.82s to 14.3s in Fig. 10(i)). Nonetheless, this analysis is performed faster in the Patch Validator because the patch causes the previously vulnerable node to exit more quickly. The bottleneck of the Symbolic DF Analysis comes from emulating memory access using MemMap. SABRE also requires an additional check for each memory write to determine whether the write corrupts $addr_{target}$ (recall Sec. 4.2). Future work could improve SABRE’s modeling of memory accesses.
@@ 6. Related Work
Control Flow Attestation. C-FLAT (Abera et al., 2016) proposed CFA by instrumenting an application binary at branch instructions with calls that trap execution in an ARM TrustZone-protected “Secure World”, where the branch destination is added to a hash chain representing the current path in a unique digest. Many techniques follow C-FLAT by combining instrumentation and TEE support to improve CFA (Armanuzzaman et al., 2025; Toffalini et al., 2019; Zhang et al., 2021; Sun et al., 2020; Wang et al., 2023; Caulfield et al., 2024a; Yadav and Ganapathy, 2023) or by employing custom hardware extensions to detect/log control flow transfers without requiring instrumentation/TEE support (Dessouky et al., 2017, 2018; Zeitouni et al., 2017; Caulfield et al., 2023).
@@
The majority of existing work in this area focuses on the correctness and security of $\sf{\mathcal{P}rv}$ implementation to generate run-time evidence (Abera et al., 2016; Toffalini et al., 2019; Zhang et al., 2021; Sun et al., 2020; Wang et al., 2023; Dessouky et al., 2017, 2018; Zeitouni et al., 2017; Nunes et al., 2021b; Caulfield et al., 2023; Yadav and Ganapathy, 2023; Neto and Nunes, 2023; Armanuzzaman et al., 2025). They often mention that with authentic evidence, $\sf{\mathcal{V}rf}$ can detect the validity of the reported path by utilizing data structures computed prior to run-time (e.g., CFG+SS (Toffalini et al., 2019; Zhang et al., 2021; Sun et al., 2020; Wang et al., 2023; Caulfield et al., 2023; Nunes et al., 2021b) or hash-sets (Dessouky et al., 2017; Zeitouni et al., 2017; Abera et al., 2016; Yadav and Ganapathy, 2023)) or by emulating execution (Dessouky et al., 2018; Nunes et al., 2021a). As demonstrated by this work, all techniques that create hash-based/hybrid evidence (E1 and E3 from Sec. 3) cannot infer the precise malicious execution path on $\sf{\mathcal{P}rv}$ and are not suitable for conducting root cause analysis.
Only two existing studies (Debes et al., 2023; Chilese et al., 2024) propose alternative approaches to CFA verification. ZEKRA (Debes et al., 2023) considers a context in which $\sf{\mathcal{V}rf}$ might also be untrusted and aims to learn secrets about $\sf{\mathcal{P}rv}$’s program through $\mathcal{C}F_{Log}$. To address this issue, ZEKRA outsources verification to a semi-dishonest worker, who uses an arithmetic circuit of CFG to generate a zero-knowledge proof of the verification process. ZEKRA’s procedures are only concerned with determining path validity, while SABRE aims to conduct further root cause analysis from $\mathcal{C}F_{Log}$. RAGE (Chilese et al., 2024) presents a technique for path verification that does not require constructing a complete CFG. RAGE trains a Graph Neural Network (GNN) on a set of $\sf{\mathcal{P}rv}$’s known $\mathcal{C}F_{Log}$-s and validates new $\mathcal{C}F_{Log}$-s using the trained GNN. Although SABRE requires constructing a partial CFG, we show this is feasible for low-end embedded applications, as apparent in Fig. 10(a). More importantly, RAGE does not support a root cause analysis from the received evidence, which is SABRE’s main goal.
Root Cause Analysis. Root-Cause Analysis (Zeller and Hildebrandt, 2002; Li et al., 2023; Yagemann et al., 2021b; Qasem et al., 2021; Liu et al., 2022; Park et al., 2024; Blazytko et al., 2020; Zeng et al., 2022; Zhou et al., 2022; Yagemann et al., 2021a) is an automated process for identifying the source of software bugs/unexpected behavior in software. Two closely related Root Cause Analysis frameworks are ARCUS and BENZENE.
ARCUS (Yagemann et al., 2021b) uses Intel PT to record an execution trace at run-time while a CFI scheme monitors execution until a control flow violation occurs. Then, it collects the execution trace and initial memory snapshot and sends them to a remote server to start emulating the malicious execution. During emulation, ARCUS performs checks for the root cause vulnerability. ARCUS and similar methods (Yagemann et al., 2021a; Zeng et al., 2022; Liu et al., 2022; Zhou et al., 2022) require a snapshot of the initial memory state. While sharing a similar goal, these prior approaches are not directly compatible with CFA. In CFA, $\sf{\mathcal{P}rv}$’s RoT may be unable to provide complete execution context, including initial memory snapshot and traces of all executing tasks since boot, particularly in on-demand attestation (Caulfield et al., 2023, 2024a; Neto and Nunes, 2023) in which $\sf{\mathcal{P}rv}$ either remains idle or executes unattested tasks in between CFA requests from $\sf{\mathcal{V}rf}$. Furthermore, storing a memory snapshot while the attested tasks execute would be infeasible for a resource-constrained $\sf{\mathcal{P}rv}$ MCU with limited memory (a common $\sf{\mathcal{P}rv}$ platform in CFA (Ammar et al., 2024)). To provide root cause analysis suitable with CFA evidence, SABRE operates without a snapshot of initial memory (which is not available from $\mathcal{C}F_{Slice}$).
BENZENE (Park et al., 2024) performs its analysis using the application binary and crash-inducing inputs. It uses a data-flow graph (DFG) to track all influenced values and mutate states to track when the program terminated successfully or crashed again. Similarly to SABRE, BENZENE starts from the crash site and performs a backtracking process. It then analyzes the crash/non-crash conditions and generates a list of likely root cause data-flow events. Unlike BENZENE, SABRE does not require data inputs and does not require constructing a DFG. Instead, SABRE combines CFG traversal with symbolic data-flow analysis to determine exploited memory instructions without requiring data inputs. Unlike BENZENE, SABRE generates and verifies a binary patch for the identified root cause.

### v3

@@ 4.1. Static Analysis and Path Verifier
$loop_{count}$ a special entry in $\mathcal{C}F_{Log}$ to optimize static loops (i.e., loops with no internal branches) (Nunes et al., 2021b; Caulfield and otehrs, 2023); denotes total times the previous $\mathcal{C}F_{Log}$ entry repeated
@@ 4.2. Exploit Locator
Phase 2. Symbolic DF Analysis. Within $\mathcal{C}F_{Slice}$ determined by Phase 1, Exploit Locator must locate the instruction that corrupts $addr_{target}$, typically requiring emulation of $\mathcal{C}F_{Slice}$ execution. Without access to run-time data prior to $\mathcal{C}F_{Slice}$ execution, Exploit Locator emulates this execution using symbolic analysis. Unlike traditional symbolic execution (King, 1976), we do not use symbolic analysis to explore different paths. Instead, we use it to reconstruct all possible data flows during $\mathcal{C}F_{Slice}$ execution (Yagemann et al., 2021a, b) and monitor for the change in data flows that result in corruption of $addr_{target}$.
@@ 4.3. Patch Generator
SABRE patching strategy (Duan et al., 2019; Xu et al., 2020; Eldefrawy et al., 2020; Zhang et al., 2017) redirects the vulnerable instructions to a secure implementation. The exact steps depend on the detected root cause (either use-after-free or buffer overflow).
4.3.1. Patching Use-After-Free
@@
4.3.2. Patching Buffer Overflows
@@
The example of Fig. 8 depicts the two steps required to obtain the buffer’s bounds. The previous phase has detected that the instruction mov r14, 0(r15) triggers a write to base, setting $addr_{Acc}$ to 0xe290. From this instruction, this phase knows that its destination register, r15, stores the write address pointing into the overflown buffer, and thus labels it as $reg_{Acc}$. To determine the start address of the overflown buffer, SABRE performs a backward traversal from 0xe290 until it finds the instruction that initializes its lower bound.
@@
Although this strategy could overestimate (i.e., the estimated bounds may include other data objects and the buffer of interest), narrowing the bounds any further cannot be done without source-code level semantics. For example, the instruction at 0xe0ea modifies the stack pointer by ten. However, it cannot be determined through the binary alone whether this is allocating space for a buffer of size ten or for all stack variables totaling ten memory addresses. Although this differentiation cannot be made through binary alone, SABRE’s approach guarantees overflows into $addr_{target}$ are blocked by restricting overflows within the buffer’s stack frame.
@@
The call instruction at 0xe106 previously called to the “unsafe” function at 0xe2fc. In the patched version, this call instruction now calls 0xe61a, the location of the patched version of the function. As depicted, this new function contains bounds-checking from T3.
@@ 4.4. Patch Validator
Figure 10. Run-time of SABRE submodules (ms) to analyze evidence of buffer overflow (OVF) and use-after-free (UAF) vulnerabilities on two MCUs and CFA architectures: hardware-based CFA atop MSP430 (Caulfield and otehrs, 2023) and TEE-based CFA atop ARM Cortex-M33 (Caulfield et al., 2024b).
@@ 5. Implementation & Evaluation
SABRE was implemented as a combination of Bash and Python and evaluated on an Ubuntu 18.04 machine with an Intel Core i7-4790 CPU at 3.60GHz and 32GB RAM. BEEBs benchmark applications for embedded platforms (Pallister et al., 2013) were used for the evaluation. They were compiled for MSP430 and ARM Cortex-M33 and executed atop unmodified CFA architectures, ACFA (Caulfield and otehrs, 2023) and TRACES (Caulfield et al., 2024b), to generate a dataset of executables and corresponding $\mathcal{C}F_{Log}$-s. The MSP430 prototype was configured with 8 KB of application program memory, and the ARM Cortex-M33 prototype was configured with 256 KB of application program memory. We believe that SABRE concepts generalize to several CFA architectures and use ACFA and TRACES as case studies due to their open-source availability.
To obtain CFGs from executables, SABRE first runs objdump over the executables to decode the assembly instructions. Then, it builds the CFG from this disassembled file. SABRE uses SymPy (Team, 2024) for evaluating symbolic expressions and Keystone assembler (Keystone, 2024) for generating ARM Cortex-M instructions. Since no MSP430 assembler exists in Python packages, we build upon and modify MSProbe (Swiftloke, 2024) for MSP430 instructions. Finally, SABRE uses the Python library ELFtools (Foundation, 2024c) to generate the final patched executable. SABRE proof of concept prototype is publicly available at (Caulfield et al., 2025).
@@ 5.1. Run-times of SABRE Sub-Modules
We evaluate the run-time of each SABRE sub-module after a malicious path was detected. Given the lack of reproducible exploits and incomplete CVEs for MCUs software (e.g., missing source codes, broken links, etc.) (Tan et al., 2024), we craft control flow attacks that exploit buffer overflow and use-after-free vulnerabilities, following a similar methodology to recent related work (Chilese et al., 2024). We then insert them into the example benchmark applications (Pallister et al., 2013) to closely mimic existing CVEs related to control flow hijacking and ROP in MCUs (National Vulnerability Database, 2021a, 2017, 2020a, 2020b, 2020c, b, 2022) to the best of our ability. For consistency across applications, we insert the vulnerability into the last function called by the main. The vulnerability is used to overwrite $addr_{target}$ to return/jump to the main function, causing an infinite loop. We execute the vulnerable programs on two unmodified CFA architectures – a TEE-based approach built atop ARM Cortex-M33 (TRACES (Caulfield et al., 2024b)) and a custom hardware-based approach built atop TI MSP430 (ACFA (Caulfield and otehrs, 2023)) – to collect $\mathcal{C}F_{Log}$-s of the attack, which are used by SABRE with the modified application binary (App.elf) for further analysis. Fig. 10 shows the run-time of each SABRE sub-module during the analysis.
@@
Path Verifier. Fig. 10(b) presents the time to verify $\mathcal{C}F_{Log}$. The verification time depends on the number of branch instructions executed by the program before the corrupted $addr_{target}$ is used. In most cases, the crafted OVF generates a larger $\mathcal{C}F_{Log}$ since it introduces more control flow transfers before corrupting $addr_{target}$ compared to the crafted UAF examples. Additionally, $\mathcal{C}F_{Log}$-s of ARM executables (generated by TRACES (Caulfield et al., 2024b)) require longer run-time to verify than $\mathcal{C}F_{Log}$-s of MSP430 executables (generated by ACFA (Caulfield and otehrs, 2023)). This is because $\mathcal{C}F_{Log}$-s from TRACES do not contain any static branch instruction destinations, whereas $\mathcal{C}F_{Log}$-s from ACFA contain all control flow transfers. Consequently, TRACES $\mathcal{C}F_{Log}$-s require more time to verify since more processing is required by $\sf{\mathcal{V}rf}$ to account for the missing entries, whereas ACFA $\mathcal{C}F_{Log}$-s can be naively followed by $\sf{\mathcal{V}rf}$ without additional effort. The detection time through $\mathcal{C}F_{Log}$ verification for the evaluated examples ranged from $\approx$1.14 to 34.74ms.
@@
Like the Exploit Locator, the run-time of the Patch Validator is dominated by the Symbolic DF Analysis (from 0.82s to 14.3s in Fig. 10(i)). Nonetheless, this analysis is performed faster in the Patch Validator because the patch causes the previously vulnerable node to exit more quickly. The bottleneck of the Symbolic DF Analysis comes from emulating memory access using MemMap. SABRE also requires an additional check for each memory write to determine whether the write corrupts $addr_{target}$ (recall Sec. 4.2). Future work could improve SABRE’s modeling of memory accesses.
@@ 6. Related Work
Control Flow Attestation. Current CFA approaches generate execution path evidence combining binary instrumentation with TEE support (Armanuzzaman et al., 2025; Toffalini et al., 2019; Zhang et al., 2021; Sun et al., 2020; Wang et al., 2023; Caulfield et al., 2024b; Yadav and Ganapathy, 2023) or using custom hardware to detect/log control flow transfers (Dessouky et al., 2017, 2018; Zeitouni et al., 2017; Caulfield and otehrs, 2023). Most CFA literature focuses on securely generating correct evidence on a compromised $\sf{\mathcal{P}rv}$, mentioning that this evidence can be used with CFG+SS (Toffalini et al., 2019; Zhang et al., 2021; Sun et al., 2020; Wang et al., 2023; Caulfield and otehrs, 2023; Nunes et al., 2021b), hash-sets (Dessouky et al., 2017; Zeitouni et al., 2017; Abera et al., 2016; Yadav and Ganapathy, 2023), or execution emulation (Dessouky et al., 2018; Nunes et al., 2021a) to perform verification. As discussed in this paper, techniques that create hash-based/hybrid evidence (E1 and E3 from Sec. 3) cannot always infer the precise malicious execution paths on $\sf{\mathcal{P}rv}$ and are not suitable for root cause analysis.
To our knowledge, only two prior efforts (Debes et al., 2023; Chilese et al., 2024) propose approaches for CFA verification. ZEKRA (Debes et al., 2023) proposes a CFA verification in zero-knowledge proofs to ensure that an untrusted $\sf{\mathcal{V}rf}$ does not learn secrets from $\mathcal{C}F_{Log}$. RAGE (Chilese et al., 2024) presents a method for path verification by training a Graph Neural Network (GNN) on $\mathcal{C}F_{Log}$-s, removing the need for a CFG. Importantly, neither ZEKRA nor RAGE supports root cause analysis from received evidence, which is SABRE’s main goal. Although SABRE requires constructing a partial CFG (unlike RAGE), we show this is feasible for MCU applications; see Fig. 10(a). Ammar et al. (Ammar et al., 2024) provide a systematization of run-time defenses and discuss CFA verification as an open challenge. Unlike their discussion on the challenges in verifying CFA evidence, our work analyzes CFA evidence formats and their suitability for advanced analysis in depth. We also show how $\sf{\mathcal{V}rf}$ can use E2 evidence to pinpoint and remediate root causes concretely.
Root Cause Analysis. Root Cause Analysis frameworks  (Yagemann et al., 2021a; Liu et al., 2022; Park et al., 2024; Zeng et al., 2022; Zhou et al., 2022; Yagemann et al., 2021b) automate the process for identifying sources of software bugs or unexpected behavior. ARCUS (Yagemann et al., 2021a) collects an Intel PT trace and an initial memory snapshot, which are used by an ARCUS server after a run-time violation to identify root cause vulnerabilities during execution emulation and to generate a report. Due to the cost of storing and transmitting large memory snapshots, ARCUS and similar proposals (Yagemann et al., 2021b; Zeng et al., 2022; Liu et al., 2022; Zhou et al., 2022) are unlikely to integrate feasibly with CFA on low-end MCUs. BENZENE (Park et al., 2024) starts with the application binary and crash-inducing inputs, and it constructs a data-flow graph (DFG) to backtrack from crashes to track state conditions that lead to crashes and non-crashes. It then analyzes these conditions to generate likely root cause data-flow events. Unlike ACRUS and BENZENE, SABRE performs root cause analysis without large memory snapshots, crash inputs, or DFGs. Instead, SABRE combines CFG traversal with symbolic data-flow analysis to identify exploited instructions, generate a binary patch for the identified root cause, and verify its effectiveness.
Binary Rewriting and Patching. Automated binary rewriting  (Wenzl et al., 2019) can be used for pre-deployment security measures (Zhang and Sekar, 2015; Zhang et al., 2013; Wang et al., 2015; Wartell et al., 2012; Kiaei et al., 2021) and for post-deployment patch generation (Duan et al., 2019; Eldefrawy et al., 2020; Zhang et al., 2017; Rajput et al., 2023; Xu et al., 2020; Williams-King et al., 2020; Duck et al., 2020). Like prior work in this space (Duan et al., 2019; Xu et al., 2020; Eldefrawy et al., 2020; Zhang et al., 2017), SABRE patch instructions redirect execution to a secure implementation. However, SABRE identifies the target through root-cause analysis instead of function matching or other binary analysis methods. Real-world deployments assume the target device has a secure software update module to install the rewritten binary. For SABRE, no additional assumptions are required since this could be implemented as the remediation function within the run-time auditing workflow (Caulfield and otehrs, 2023; Caulfield et al., 2024b). Deployment challenges, like vendor-specific limits placed on software updates or flash endurance levels, would be accounted for in implementing the remediation function, and thus they are outside the scope of prior works in binary rewriting and SABRE.